Shift report 20020101 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Jan 01 2002 - 00:13:20 EST

  • Next message: Apache: "Shift report 20020101 08:00-16:00"

    ------------------------------------------------------------
    Date: Tue Jan  1 00:13:20 2002, Shift: 16:00-24:00
    Supervisor: sjs, with: 
    Log book#11 pages 128 - 129
    
    Shift Summary:
    ------------------------------------------------------------
    Beam stored and cogged since 16:50
    
    Runs 6227 (69.7K trig 6), 6228(34K), 6229(40K)
    
    NOTE: The FS angle in DB incorrectly entered as 8 deg. since 
    run 6220. This should be changed!  The correct  FS 
    angle for run 6220 through 6229, inclusive, is 12 deg.
    
    We have accumlated a total of ~200000 trig 6 events
    with this FS setting (12 deg.!!). 
    The goal is for 300000 events.
    
    At the beginning of the shift, before run 6227, the 
    attenuation and DC offsets were adjusted for si1_1c to
    see if we could move the single mip peak away from the 
    pedestal.  The experiment (done with HI) indicates that 
    it might be necessary to change some of the resistor values 
    of the attenuator circuit to obtain the desired result of
    high gain and correct dc offset...
    
    HAPPY NEW YEAR!!!  (It just passed midnight...)
    
    Only potentially new detector problem is that for the 
    tofwcal hit dist (Top+bot) spectrum, slot 2 has a low number of hits
    compared to earlier runs. 
    
    After I submit this report I plan to call it quits and
    inform MCR that we will be down for about 9 hours....The
    day shift may start a little late tomorrow...
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Jan 01 2002 - 00:13:54 EST